# Utilizing the Janus MoSSe surface polarization property in complementary metal–oxide–semiconductor field effect transistor design

### v2

@@ III Results and discussion
The – characteristics of several MoSSe MOSFET designs are now investigated. Figure 5(a) shows a single-gate design with gate length m, which is a common dimension adopted in experimental studies of FETs based on 2D materialsXiong et al. (2021); Sebastian et al. (2021). Device modifications such as the use of double gates [Fig. 5(b)] and smaller dimension [Fig. 5(c)] are also simulated. The simulations for all devices assume an HfO2 gate oxide of thickness nm. Figure 5(d) shows the transfer characteristics (–) of the n- and p-MOSFETs from Figs. 5(a)–5(c). The curves with squares, circles, and diamonds correspond to the n-MOSFETs in (a), (b), and (c), respectively, and those with triangles, stars, and pentagons to the p-MOSFETs in (a), (b), and (c), respectively. The key distinction lies in the opposite orientation of in the JTMD layer. The typical – curves for the n- and p-MOSFETs indicate that all three devices exhibit satisfactory enhancement-mode MOSFET behavior. Among the three devices, the single-gate design [Figure 5(a)] has the most significant sub-threshold swing (SS) and the lowest on-current . When a bottom gate is added [Figure 5(b)], the SS drops to around 60.5 mV/dec, while increases by an order of magnitude, indicating an improvement in gate control. Having shown that both designs are feasible at a larger scale, we reduce the horizontal scale of the double-gate design to nm [Figure 5(c)] to examine the device performance at a smaller scale. The proximity of the SS values in the n-MOSFET suggests that reducing to 200 nm would not significantly impact the control of the gate in this design. In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the off-state current is decreased by about an order, indicating an improvement in . All of these results show that the SS and on/off-current characteristics of the designs considered are those of a typical MOSFET device. The lower SS and higher consistently observed for all the n-MOSFETs compared with the corresponding p-MOSFETs may originate from the slightly higher mobility of electrons compared with holes.
@@
For the successful operation of a CMOSFET, an enhancement-mode MOSFET device is essential. Since the gate oxide thickness plays a crucial role in determining the operation mode of a MOSFET, a variety of oxides and different values of , are tested to determine the optimal design of the device in Fig. 6(a) in terms of the equivalent oxide thickness (EOT), and the results are shown in Figs. 6(b) and 6(c) for the n- and p-MOSFETs, respectively. Here, we define on/off state switching when . First, a SiO2 oxide gate of thickness 1 nm is tested, and the – result is shown by the curve with red circles. The device is clearly in the on-state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device. To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an Al2O3 gate with nm (EOT 0.42 nm, blue triangles). However, the current remains of the order of A/cm at V, and both polarities remain in the depletion mode. Therefore, a high- material, namely, HfO2, is selected to help decrease the EOT of the MOSFETs. Using HfO2 with nm (EOT 0.16 nm, green stars) gives a superior enhancement-mode device for both n- and p-MOSFETs. Finally, we optimize the of the HfO2 to 2 nm (EOT 0.31 nm, purple diamonds) to prevent current leakage through the thin gate oxide. As can be seen from Figs. 6(b) and 6(c), both of the – curves for this optimized value of intersect the off-state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased. These results for HfO2 with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm. For this optimal MOSFET design based on HfO2 with nm, the SS as extracted from Figs. 6(b) and 6(c) is about 63 mV/dec. Since such low- devices may encounter several processing difficulties, simulation results for a more easily fabricated thick HfO2 gate with nm (EOT 1.6 nm, black squares) are also given in Figs. 6(b) and 6(c). Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 2.1 V and 1.8 V for n- and p-MOSFETs, respectively. All in all, the – plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
@@
It is also worth considering how the value of affects this carrier accumulation phenomenon at the channel MoSSe region. To reveal the impact of the intrinsic dipole’s magnitude on device performance, we calculate the two-dimensional electron gas (2DEG) in the n-MOSFET device with a 200 nm length channel, shown in [Figure 7(a)], with various assumed values ranging from cm-2 to cm-2. These values can be regarded as JTMDs with different intrinsic dipole strengths, providing insights into how the magnitude influences the device’s 2DEG. The 2DEG corresponding to MoSSe’s is specified in the black dot frame. SiO2 ( and nm) and HfO2 ( and nm) were selected as representative low- and high- gate oxides for this study. To isolate the impact of , the calculations were performed at zero bias, corresponding to the equilibrium state.
@@
[Figure 7(b)] illustrates the influence of on the 2DEG concentration. In cases with high EOT (SiO2, nm), 2DEG exhibits a nearly linear increase with increasing . This linearity arises because the high EOT allows carriers to reach saturation within the investigated range ( to cm-2). However, as EOT decreases, the starting point of the linear region shifts towards higher values. For instance, the linear region begins at cm-2 for HfO2 with nm and cm-2 for SiO2 with nm. The value at the onset of the linear regime serves as an indicator for achieving enhanced mode devices. The HfO2 design with nm has the lowest EOT, resulting in a linear regime starting beyond cm-2. This suggests that MoSSe (marked by the brown dotted box) will operate in enhanced mode due to the initially low 2DEG concentration. Conversely, increasing the HfO2 thickness to 3 nm lowers the EOT and brings the linear region starting point to cm-2. This indicates a depletion mode behavior for MoSSe with HfO2 ( nm) due to the higher initial induced 2DEG.
@@
The influence of extends to the IV characteristics of MOSFETs. To focus on its impact on the linear 2DEG region, we examined devices with values of , , and cm-2, all designed with a 10 nm thick HfO2. As shown in Figure 7(c), an increased shifts the transfer IV curve towards more negative voltages. This shift reflects the need for a stronger gate voltage to deplete the augmented 2DEG in the channel. Moreover, the higher carrier concentration resulting from a higher leads to a greater on-current (Ion). Additionally, can influence carrier injection at the contacts. For strengths of and cm-2 value of , the device also has a larger gm value due to a smaller contact resistance. However, if falls below a certain threshold, devices display non-ideal IV curves near saturation (highlighted by the purple dotted box). As detailed in Figure 4(c), insufficient compromises Ohmic contacts, resulting in increased resistance from an additional barrier at the contact regions. This issue can be mitigated by selecting a metal with a work function closer to the band edge to offset the weaker band bending at the contact interface. For p-MOSFETs, the same principles apply.
@@
To understand the distribution of carriers in the device, we export the two-dimensional distribution map of the majority carrier of the device under different operating conditions in Figure 8, which demonstrates the on/off characteristics of the p- and n-MOSFET devices with HfO2 gate with nm. The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm-3 (green color). The carrier concentration gradually decays as the position moves away from the contact region. Owing to the 10 nm gate metal isolation design, two bump-like regions of high carrier concentration appear next to the source and drain areas. The thicker oxide eliminates the dipole effect of MoSSe, and thus carrier gas remains. Taking the p-MOSFET as an example, we can see that with the application of a negative 0.5 V gate bias in Fig. 8(a), holes are attracted into the channel region, forming a strong inversion channel layer and causing the device to turn on. By contrast, when a positive 0.5 V gate bias is applied in Fig. 8(b), holes are depleted, and the inversion layer disappears, indicating an off-state device. A similar result is also obtained for the n-MOSFET, with the only difference being that a strong inversion layer appears when V (Fig. 8(c)) and the inversion layer vanishes when a V bias is applied (Fig. 8(d)).
@@ IV Conclusions
We have introduced a novel CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe. Our investigation has encompassed diverse heterostructures, including Gr/AA′A-MoSSe/Gr and 2L-MoSSe/Al (111) face and 2L-MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT. We have also examined the interactions between MoSSe, Gr, and Al and Cu electrodes, identifying characteristics such as Schottky barriers and Ohmic contacts. Our design leverages the strong dipole field of MoSSe to create n- and p-type contact regions, thereby forming a CMOSFET with simplified material selection. The proposed horizontal CMOSFET architecture offers advantages in fabrication reliability and integration. Device simulations have provided insights into the spontaneous polarization caused by band bending and the related carrier distribution, with an optimized EOT 0.312 nm giving a favorable threshold between the depletion and enhancement modes of the MOSFET device. The discussion also delves into the influence of spontaneous polarization magnitude, a crucial factor determining the conditions for achieving Ohmic contact and the formation of the channel 2DEG/2DHG. The majority of carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits. On the basis of the surface polarization strategy adopted here, it might be possible to form a MOSFET from any JTMD, provided that the 2D material possesses a strong intrinsic out-of-plane dipole. Regardless, this work has introduced a promising approach to CMOSFET design, offering valuable insights into materials, interfaces, and device characteristics for future electronic applications.
@@ Acknowledgements.
This work was supported by the National Science and Technology Council under Grant Nos. NSTC 112-2221-E-002-215-MY3, NSTC 112-2923-E-002-002, and 12-2221-E-002-214-MY3.AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose
Author Contributions
Yun-Pin Chiu: Conceptualization (equal); Data curation (lead);Formal analysis (equal); Methodology (equal); Validation (lead); Visualization(lead); Writing – original draft (lead). Hsin-Wen Huang: Conceptualization (equal); Data curation (support); Formal analysis (equal); Methodology (equal); Validation (support). Yuh-Renn Wu: Conceptualization (equal); Formal analysis (equal); Funding acquisition (lead); Project administration (lead); Software(lead); Supervision (lead); Writing – review and editing (lead).
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v3

@@ III Results and discussion
The - characteristics of several MoSSe MOSFET designs are now investigated. Figure 5(a) shows a single-gate design with gate length m, which is a common dimension adopted in experimental studies of FETs based on 2D materialsXiong et al. (2021); Sebastian et al. (2021). Device modifications such as the use of double gates [Fig. 5(b)] and smaller dimension [Fig. 5(c)] are also simulated. The simulations for all devices assume an HfO2 gate oxide of thickness nm. Figure 5(d) shows the transfer characteristics (-) of the n- and p-MOSFETs from Figs. 5(a)–5(c). The curves with squares, circles, and diamonds correspond to the n-MOSFETs in (a), (b), and (c), respectively, and those with triangles, stars, and pentagons to the p-MOSFETs in (a), (b), and (c), respectively. The key distinction lies in the opposite orientation of in the JTMD layer. The typical - curves for the n- and p-MOSFETs indicate that all three devices exhibit satisfactory enhancement-mode MOSFET behavior. Among the three devices, the single-gate design [Figure 5(a)] has the most significant subthreshold swing (SS) and the lowest ONcurrent . When a bottom gate is added [Figure 5(b)], the SS drops to around 60.5 mV/dec, while increases by an order of magnitude, indicating an improvement in gate control. Having shown that both designs are feasible at a larger scale, we reduce the horizontal scale of the double-gate design to nm [Figure 5(c)] to examine the device performance at a smaller scale. The proximity of the SS values in the n-MOSFET suggests that reducing to 200 nm would not significantly impact the control of the gate in this design. In addition, is increased by an order of magnitude compared with that of the larger-scale device, while the OFF-state current is decreased by about an order, indicating an improvement in . All of these results show that the SS and ON- or OFF-current characteristics of the designs considered are those of a typical MOSFET device. The lower SS and higher consistently observed for all the n-MOSFETs compared with the corresponding p-MOSFETs may originate from the slightly higher mobility of electrons compared with holes.
@@
For the successful operation of a CMOSFET, an enhancement-mode MOSFET device is essential. Since the gate oxide thickness plays a crucial role in determining the operation mode of a MOSFET, a variety of oxides and different values of , are tested to determine the optimal design of the device in Fig. 6(a) in terms of the equivalent oxide thickness (EOT), and the results are shown in Figs. 6(b) and 6(c) for the n- and p-MOSFETs, respectively. Here, we define ON or OFFstate switching when . First, a SiO2 oxide gate of thickness 1 nm is tested, and the - result is shown by the curve with red circles. The device is clearly in the ON state ( A/cm) when no gate bias is applied, indicating that the MOSFET is a depletion-mode device. To search for the threshold EOT at which the device switches to the enhancement mode, we try decreasing the EOT by selecting an Al2O3 gate with nm (EOT 0.42 nm, blue triangles). However, the current remains of the order of A/cm at V, and both polarities remain in the depletion mode. Therefore, a high- material, namely, HfO2, is selected to help decrease the EOT of the MOSFETs. Using HfO2 with nm (EOT 0.16 nm, green stars) gives a superior enhancement-mode device for both n- and p-MOSFETs. Finally, we optimize the of the HfO2 to 2 nm (EOT 0.31 nm, purple diamonds) to prevent current leakage through the thin gate oxide. As can be seen from Figs. 6(b) and 6(c), both of the – curves for this optimized value of intersect the OFF state brown dashed line at V, indicating preservation of the enhancement mode even though the EOT has increased. These results for HfO2 with nm suggest that the EOT threshold between the depletion and enhancement modes in this MOSFET design is 0.312 nm. For this optimal MOSFET design based on HfO2 with nm, the SS as extracted from Figs. 6(b) and 6(c) is about 63 mV/dec. Since such low- devices may encounter several processing difficulties, simulation results for a more easily fabricated thick HfO2 gate with nm (EOT 1.6 nm, black squares) are also given in Figs. 6(b) and 6(c). Owing to the increased EOT, both the n- and p-type MOSFETs exhibit the depletion mode in terms of transfer characteristics, where both the drops and the device turns off at 2.1 V and 1.8 V for n- and p-MOSFETs, respectively. All in all, the - plots demonstrate the possibility of creating a workable MOSFET even with larger values of and provide a reference for future experimental studies.
@@
It is also worth considering how the value of affects this carrier accumulation phenomenon at the channel MoSSe region. To reveal the impact of the intrinsic dipole’s magnitude on device performance, we calculate the two-dimensional electron gas (2DEG) in the n-MOSFET device with a 200-nm length channel, shown in [Figure 7(a)], with various assumed values ranging from to cm-2. These values can be regarded as JTMDs with different intrinsic dipole strengths, providing insights into how the magnitude influences the device’s 2DEG. The 2DEG corresponding to MoSSe’s is specified in the black dot frame. SiO2 ( and nm) and HfO2 ( and nm) were selected as representative low- and high- gate oxides for this study. To isolate the impact of , the calculations were performed at zero bias, corresponding to the equilibrium state.
@@
Figure 7(b) illustrates the influence of on the 2DEG concentration. In cases with high EOT (SiO2, nm), 2DEG exhibits a nearly linear increase with increasing . This linearity arises because the high EOT allows carriers to reach saturation within the investigated range ( to cm-2). However, as EOT decreases, the starting point of the linear region shifts towards higher values. For instance, the linear region begins at cm-2 for HfO2 with nm and cm-2 for SiO2 with nm. The value at the onset of the linear regime serves as an indicator for achieving enhanced mode devices. The HfO2 design with nm has the lowest EOT, resulting in a linear regime starting beyond cm-2. This suggests that MoSSe (marked by the brown dotted box) will operate in enhanced mode due to the initially low 2DEG concentration. Conversely, increasing the HfO2 thickness to 3 nm lowers the EOT and brings the linear region starting point to cm-2. This indicates a depletion mode behavior for MoSSe with HfO2 ( nm) due to the higher initial induced 2DEG.
@@
The influence of extends to the I-V characteristics of MOSFETs. To focus on its impact on the linear 2DEG region, we examined devices with values of , , and cm-2, all designed with a 10 nm thick HfO2. As shown in Figure 7(c), an increased shifts the transfer I-V curve towards more negative voltages. This shift reflects the need for a stronger gate voltage to deplete the augmented 2DEG in the channel. Moreover, the higher carrier concentration resulting from a higher leads to a greater ON current (ION). Additionally, can influence carrier injection at the contacts. For strengths of and cm-2 value of , the device also has a larger gm value due to a smaller contact resistance. However, if falls below a certain threshold, devices display nonideal I-V curves near saturation (highlighted by the purple dotted box). As detailed in Figure 4(c), insufficient compromises Ohmic contacts, resulting in increased resistance from an additional barrier at the contact regions. This issue can be mitigated by selecting a metal with a work function closer to the band edge to offset the weaker band bending at the contact interface. For p-MOSFETs, the same principles apply.
@@
To understand the distribution of carriers in the device, we export the two-dimensional distribution map of the majority carrier of the device under different operating conditions in Figure 8, which demonstrates the ONorOFF characteristics of the p- and n-MOSFET devices with HfO2 gate with nm. The source and drain areas are defined as those where the MoSSe is attached to the metal, which gives a relatively high carrier concentration of about cm-3 (green color). The carrier concentration gradually decays as the position moves away from the contact region. Owing to the 10-nm gate metal isolation design, two bumplike regions of high carrier concentration appear next to the source and drain areas. The thicker oxide eliminates the dipole effect of MoSSe, and thus carrier gas remains. Taking the p-MOSFET as an example, we can see that with the application of a negative 0.5 V gate bias in Fig. 8(a), holes are attracted into the channel region, forming a strong inversion channel layer and causing the device to turn on. By contrast, when a positive 0.5 V gate bias is applied in Fig. 8(b), holes are depleted, and the inversion layer disappears, indicating an OFF-state device. A similar result is also obtained for the n-MOSFET, with the only difference being that a strong inversion layer appears when V (Fig. 8(c)) and the inversion layer vanishes when a V bias is applied (Fig. 8(d)).
@@ IV Conclusions
We have introduced an alternative CMOSFET design based on the intrinsic dipole properties of JTMD, with a specific focus on the material MoSSe. Our investigation has encompassed diverse heterostructures, including Gr/AA′A-MoSSe/Gr and 2L-MoSSe/Al (111) face and 2L-MoSSe/Cu (111) face structures, which we have analyzed using first-principles DFT. We have also examined the interactions between MoSSe, Gr, and Al and Cu electrodes, identifying characteristics, such as Schottky barriers and Ohmic contacts. Our design leverages the strong dipole field of MoSSe to create n- and p-type contact regions, thereby forming a CMOSFET with simplified material selection. The proposed horizontal CMOSFET architecture offers advantages in fabrication reliability and integration. Device simulations have provided insights into the spontaneous polarization caused by band bending and the related carrier distribution, with an optimized EOT 0.312 nm giving a favorable threshold between the depletion and enhancement modes of the MOSFET device. The discussion also delves into the influence of spontaneous polarization magnitude, a crucial factor determining the conditions for achieving Ohmic contact and the formation of the channel 2DEG and 2DHG. The majority of carrier plots show the capability of the device to control carriers in the MoSSe channel region, demonstrating its potential for low-voltage logic elements in integrated circuits. On the basis of the surface polarization strategy adopted here, it might be possible to form a MOSFET from any JTMD, provided that the 2D material possesses a strong intrinsic out-of-plane dipole. Regardless, this work has introduced a promising approach to CMOSFET design, offering valuable insights into materials, interfaces, and device characteristics for future electronic applications.
@@ Acknowledgements.
This work was supported by the National Science and Technology Council under Grants No. NSTC 112-2221-E-002-215-MY3, No. NSTC 112-2923-E-002-002, and No. 112-2221-E-002-214-MY3. Yun-Pin Chiu: conceptualization (equal); data curation (lead); formal analysis (equal); methodology (equal); validation (lead); visualization(lead); writing – original draft (lead). Hsin-Wen Huang: conceptualization (equal); data curation (support); formal analysis (equal); methodology (equal); validation (support). Yuh-Renn Wu conceptualization (equal); formal analysis (equal); funding acquisition (lead); project administration (lead); software(lead); supervision (lead); writing-–review and editing (lead). The authors have no conflicts to discloseReferences